# ROSE: Register Assisted General Time Series Forecasting with Decomposed Frequency Learning

### v1

@@ H.3 Ablation study results
avg 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.38
Replace Multi Freq Masking Single Freq Masking 96 0.296 0.346 0.163 0.253 0.371 0.398 0.272 0.335
192 0.336 0.368 0.221 0.291 0.406 0.420 0.340 0.376
336 0.362 0.385 0.276 0.327 0.411 0.426 0.379 0.406
720 0.424 0.418 0.366 0.381 0.422 0.447 0.399 0.431
avg 0.354 0.379 0.257 0.313 0.403 0.423 0.347 0.387
Multi Patch Masking 96 0.302 0.348 0.168 0.257 0.377 0.408 0.282 0.343
@@
H.4 Zero-shot results
Table 11: Full results in zero-shot setting
Model ROSE MOIRAI Timer PatchTST GPT4TS
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.512 0.460 0.660 0.476 0.596 0.519 0.906 0.578 0.743 0.545
192 0.512 0.462 0.707 0.500 0.668 0.560 0.892 0.573 0.759 0.563
336 0.523 0.470 0.730 0.515 1.580 0.871 0.876 0.567 0.748 0.570
720 0.552 0.490 0.758 0.536 1.587 0.882 0.836 0.554 0.780 0.576
avg 0.525 0.471 0.714 0.507 1.108 0.708 0.877 0.568 0.758 0.564
ETTm2 96 0.224 0.309 0.216 0.282 0.221 0.300 0.301 0.346 0.275 0.343
192 0.266 0.333 0.294 0.330 0.303 0.349 0.337 0.367 0.298 0.356
336 0.310 0.358 0.368 0.373 0.411 0.423 0.376 0.389 0.347 0.386
720 0.395 0.407 0.494 0.439 0.505 0.473 0.456 0.434 0.426 0.417
avg 0.299 0.352 0.343 0.356 0.360 0.386 0.367 0.384 0.337 0.376
ETTh1 96 0.382 0.408 0.405 0.397 0.414 0.439 0.583 0.502 0.573 0.521
192 0.400 0.420 0.458 0.428 0.458 0.465 0.600 0.516 0.594 0.529
336 0.404 0.426 0.509 0.454 0.475 0.479 0.594 0.525 0.682 0.567
720 0.420 0.447 0.529 0.494 0.521 0.511 0.569 0.544 0.543 0.524
avg 0.401 0.425 0.475 0.443 0.467 0.474 0.586 0.522 0.598 0.535
ETTh2 96 0.298 0.362 0.303 0.338 0.313 0.366 0.401 0.418 0.369 0.411
192 0.336 0.385 0.369 0.384 0.394 0.415 0.412 0.425 0.382 0.423
336 0.353 0.399 0.397 0.410 0.420 0.439 0.410 0.427 0.401 0.442
720 0.395 0.432 0.447 0.450 0.440 0.459 0.415 0.437 0.424 0.439
avg 0.346 0.394 0.379 0.396 0.392 0.420 0.409 0.427 0.394 0.429

### v2

@@ A.10.3 Ablation study results
avg 0.349 0.372 0.250 0.308 0.397 0.419 0.335 0.380
Replace Multi-Frequency Masking Random Frequency Masking 96 0.330 0.370 0.170 0.262 0.391 0.399 0.303 0.359
192 0.352 0.392 0.232 0.298 0.406 0.430 0.356 0.395
336 0.390 0.389 0.276 0.342 0.411 0.432 0.417 0.428
720 0.452 0.438 0.366 0.392 0.432 0.447 0.420 0.439
avg 0.381 0.397 0.261 0.324 0.410 0.427 0.374 0.405
Multi-Patch Masking 96 0.302 0.348 0.168 0.257 0.377 0.408 0.282 0.343
@@
From Scratch 96 0.301 0.357 0.171 0.260 0.419 0.439 0.315 0.389
192 0.358 0.385 0.223 0.294 0.438 0.457 0.366 0.405
336 0.390 0.396 0.282 0.336 0.484 0.484 0.424 0.435
720 0.436 0.427 0.366 0.380 0.540 0.538 0.495 0.473
avg 0.371 0.391 0.261 0.318 0.470 0.480 0.400 0.425
A.10.4 Zero-shot results
Table 15: Full results in zero-shot setting.
Models ROSE_512 Timer MOIRAI Chronos TimesFM Moment
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.382 0.408 0.414 0.439 0.405 0.397 0.494 0.409 0.432 0.405 0.706 0.561
192 0.400 0.420 0.440 0.455 0.458 0.428 0.561 0.443 0.492 0.438 0.716 0.579
336 0.404 0.426 0.455 0.463 0.509 0.454 0.580 0.460 0.519 0.458 0.705 0.583
720 0.420 0.447 0.496 0.496 0.529 0.494 0.605 0.495 0.512 0.477 0.705 0.597
avg 0.401 0.425 0.451 0.463 0.475 0.443 0.560 0.452 0.489 0.444 0.708 0.580
ETTh2 96 0.298 0.362 0.305 0.355 0.303 0.338 0.306 0.338 0.311 0.345 0.373 0.416
192 0.336 0.385 0.365 0.406 0.369 0.384 0.396 0.394 0.401 0.397 0.384 0.422
336 0.353 0.399 0.378 0.413 0.397 0.410 0.423 0.417 0.436 0.430 0.386 0.426
720 0.395 0.432 0.414 0.457 0.447 0.450 0.442 0.439 0.437 0.450 0.425 0.454
avg 0.346 0.394 0.366 0.408 0.379 0.396 0.392 0.397 0.396 0.405 0.392 0.430
ETTm1 96 0.512 0.460 0.440 0.422 0.660 0.476 0.514 0.443 0.366 0.374 0.679 0.544
192 0.512 0.462 0.505 0.458 0.707 0.500 0.608 0.475 0.413 0.401 0.690 0.550
336 0.523 0.470 0.570 0.490 0.730 0.515 0.690 0.507 0.445 0.429 0.701 0.557
720 0.552 0.490 0.659 0.534 0.758 0.536 0.733 0.555 0.513 0.470 0.719 0.569
avg 0.525 0.471 0.544 0.476 0.714 0.507 0.636 0.495 0.434 0.419 0.697 0.555
ETTm2 96 0.224 0.309 0.203 0.285 0.216 0.282 0.202 0.293 0.189 0.257 0.230 0.308
192 0.266 0.333 0.265 0.327 0.294 0.330 0.286 0.348 0.277 0.325 0.285 0.338
336 0.310 0.358 0.319 0.361 0.368 0.373 0.355 0.386 0.350 0.381 0.339 0.369
720 0.395 0.407 0.405 0.410 0.494 0.439 0.409 0.425 0.464 0.448 0.423 0.424
avg 0.299 0.352 0.360 0.386 0.343 0.356 0.313 0.363 0.320 0.353 0.319 0.360
Weather 96 0.200 0.260 0.190 0.236 0.188 0.250 0.209 0.244 - - 0.216 0.271
192 0.239 0.288 0.261 0.293 0.237 0.284 0.254 0.288 - - 0.264 0.306
336 0.279 0.315 0.332 0.340 0.282 0.323 0.301 0.332 - - 0.313 0.336
720 0.340 0.357 0.385 0.381 0.359 0.345 0.388 0.374 - - 0.369 0.380
avg 0.265 0.305 0.292 0.312 0.267 0.300 0.288 0.310 - - 0.291 0.323
Electricity 96 0.209 0.307 0.210 0.312 0.212 0.301 0.194 0.266 - - 0.844 0.761
192 0.219 0.315 0.239 0.337 0.225 0.320 0.218 0.289 - - 0.850 0.762
336 0.236 0.330 0.284 0.372 0.245 0.333 0.244 0.321 - - 0.862 0.766
720 0.273 0.328 0.456 0.479 0.282 0.358 0.324 0.371 - - 0.888 0.774
avg 0.234 0.320 0.297 0.375 0.241 0.328 0.245 0.312 - - 0.861 0.766
Traffic 96 0.572 0.407 0.526 0.368 - - 0.562 0.378 - - 1.390 0.800
192 0.575 0.406 0.561 0.385 - - 0.579 0.412 - - 1.403 0.802
336 0.588 0.411 0.614 0.412 - - 0.594 0.420 - - 1.415 0.804
720 0.618 0.422 0.749 0.464 - - 0.723 0.472 - - 1.437 0.808
avg 0.588 0.412 0.613 0.407 - - 0.615 0.421 - - 1.411 0.804